OpenStudy (blaque678):

The sum of two numbers is 42 and the difference is 12.What are the numbers?

OpenStudy (cwtan):

x+y=42 x-y=12 x=12+y 12+y+y=42 12+2y=42 2y=30 y=15 x=12+15=27

OpenStudy (cwtan):

(Assume x and y are the numbers)

OpenStudy (anonymous):

x + (x+12)=42 2x +12 =42 2x =30 x=15

OpenStudy (blaque678):

It should be 2 numbers.One large number and one small number.